OpenStudy (anonymous):

sec theta+tan theta-1/ tan theta - sec theta +1=cos theta/1- sin theta

OpenStudy (anonymous):

\[\frac{ \sec \theta+\tan \theta-1 }{ \ \tan \theta -\ \sec \theta +1 }=\frac{ \cos \theta }{ 1-\sin \theta }\] L.H.S=\[\frac{ \sec \theta +\tan \theta-1 }{ \tan \theta -\sec \theta+1 }=\frac{ \sec \theta +\tan \theta -\left( \sec ^2\theta-\tan ^2\theta \right) }{ \tan \theta-\sec \theta +1 }\] \[=\frac{ \sec \theta+\tan \theta-\left( \sec \theta +\tan \theta \right)\left( \sec \theta -\tan \theta \right) }{ \tan \theta-\sec \theta +1}\] \[=\frac{ \left( \sec \theta +\tan \theta \right)\left( 1-\sec \theta+\tan \theta \right) }{ \tan \theta -\sec \theta+1 }\] \[=(\sec \theta+\tan \theta)*\frac{ \sec \theta -\tan \theta }{ \sec \theta -\tan \theta }=\frac{ \sec ^2\theta -\tan ^2 \theta}{ \sec \theta -\tan \theta }\] \[=\frac{ 1 }{ \sec \theta -\tan \theta }=?\]